2x+7y=8
5x-3y=21
por los tres metodos please

Answer by mananth(16949)

2x+7y=8..........1
5x-3y=21 ........2
..
multiply equation 1 by -5 and equation 2 by 2 and add them
-10x-35y+10x-6y=-40+42
-41y=+2
y=-2/41
..
plug value of x in equation 1
2x +7(-2/41)= 8
2x= 8+(14/41)
2x=328+14/41
divide by 2
y= 342/ 41*2
y=171/7